### v2

@@ C Demand Estimation
Before discussing each set of IVs, we present our identification strategy. Our approach builds on a key insight from the literature that the airlines’ decisions to establish a presence in some markets are long-term strategic choices that are unlikely to depend on short-term demand or cost shocks. If so, we can construct IVs based on carriers’ network structures. In particular, for each market, we create a vector of IVs that measure the number of markets each major airline serves from the origin and destination airports. This approach captures the scope of carriers’ networks, which are largely fixed in the short run but affect pricing decisions and the provision of nonstop service.
@@
Note: The left panel shows price coefficients, and the right panel displays the kernel density estimates of implied unobserved product characteristics $\xi_{j}$ from demand estimation for all three mergers (DL-NW, UA-CO, and AA-US). For each merger, we estimate demand using seven specifications: (1-2) markets connecting top 50 MSAs three years pre-merger; (3) markets connecting top 50 MSAs three years post-merger; (4) markets connecting top 50 MSAs one year post-merger; (5) three-year post-merger sample with exogenous variables from three years pre-merger; (6) one-year post-merger sample with exogenous variables from one-year pre-merger; and (7) three-year post-merger sample with exogenous variables from one-year pre-merger. These specifications allow us to examine the stability of demand parameters across different time horizons and test the exogeneity of our instruments.

### v3

@@ C Demand Estimation
Before discussing each set of IVs, we present our identification strategy. Our approach builds on a key insight from the literature that the airlines’ decisions to establish a presence in some markets are long-term strategic choices unlikely to depend on short-term demand or cost shocks. If so, we can construct IVs based on carriers’ network structures. In particular, for each market, we create a vector of IVs that measure the number of markets each major airline serves from the origin and destination airports. This approach captures the scope of carriers’ networks, which are largely fixed in the short run but affect pricing decisions and the provision of nonstop service.
@@
Note: The left panel shows price coefficients and the right panel displays the kernel density estimates of implied unobserved product characteristics $\xi_{j}$ from demand estimation for all three mergers (DL-NW, UA-CO, and AA-US). For each merger, we estimate demand using seven specifications: (1-2) markets connecting top 50 MSAs three years pre-merger; (3) markets connecting top 50 MSAs three years post-merger; (4) markets connecting top 50 MSAs one-year post-merger; (5) three-year post-merger sample with exogenous variables from three years pre-merger; (6) one-year post-merger sample with exogenous variables from one-year pre-merger; and (7) three-year post-merger sample with exogenous variables from one-year pre-merger. These specifications allow us to examine the stability of demand parameters across different time horizons and test the exogeneity of our instruments.